# Enhancing Fourier Neural Operators with Local Spatial Features

### v1

@@ Abstract
Partial Differential Equation (PDE) problems often exhibit strong local spatial structures, and effectively capturing these structures is critical for approximating their solutions. Recently, the Fourier Neural Operator (FNO) has emerged as an efficient approach for solving these PDE problems. By using parametrization in the frequency domain, FNOs can efficiently capture global patterns. However, this approach inherently overlooks the critical role of local spatial features, as frequency-domain parameterized convolutions primarily emphasize global interactions without encoding comprehensive localized spatial dependencies. Although several studies have attempted to address this limitation, their extracted Local Spatial Features (LSFs) remain insufficient, and computational efficiency is often compromised. To address this limitation, we introduce a convolutional neural network (CNN) preprocessor to extract LSFs directly from input data, resulting in a hybrid architecture termed Conv-FNO. Furthermore, we introduce two novel resizing schemes to make our Conv-FNO resolution invariant. In this work, we focus on demonstrating the effectiveness of incorporating LSFs into FNOs by conducting both a theoretical analysis and extensive numerical experiments. Our findings show that this simple yet impactful modification enhances the representational capacity of FNOs and significantly improves performance on challenging PDE benchmarks.
@@ 1 Introduction
Inspired by the success of CNO, we propose a simple yet powerful enhancement to the FNO architecture: the Convolutional Fourier Neural Operator (Conv-FNO). Specifically, our approach integrates a convolutional neural network (CNN) as a preprocessor to extract LSFs directly from the input data before applying the Fourier transform. This preprocessing step effectively bridges the gap between local and global representations, allowing the model to retain fine-grained spatial details while leveraging the Fourier transform for efficient global interactions. Unlike previous methods focusing on frequency-domain modifications or downstream architectural enhancements, our approach explicitly targets input-stage feature extraction to ensure that critical local information is preserved from the outset.
Our proposed Conv-FNO preserves the resolution-invariance property of the original FNO through carefully designed resizing schemes, making it adaptable to grids of varying resolutions. Additionally, we introduce a specific variant, UNet-FNO, which incorporates a UNet-based encoder-decoder structure [26, 19] as the CNN preprocessor. This design leverages the hierarchical feature extraction capabilities of UNet to capture multi-scale spatial information, significantly enhancing the model’s ability to handle complex PDE solutions requiring both local and global representations.
@@ 4.2 Theoretical Analysis
The goal of an FNO is to approximate an operator between two Hilbert spaces $\mathcal{H}_{1},\mathcal{H}_{2}$, which we call $\mathcal{C}:\mathcal{H}_{1}\to\mathcal{H}_{2}$. We now suppose that the Hilbert space $\mathcal{H}_{1}$ is made of functions from $\Omega\subset\mathcal{R}^{n}$ and taking outputs in $\mathcal{R}^{d_{1}}$. The Hilbert space $\mathcal{H}_{2}$ takes instead inputs in ${\Omega}\subset\mathcal{R}^{n}$ and returns outputs in $\mathcal{R}^{d_{2}}$. This means that for an $x\in\Omega$, $u\in\mathcal{H}_{1}$, we have $u(x)\in\mathcal{R}^{d_{1}}$ and $\mathcal{C}(u)(x)\in\mathcal{R}^{d_{2}}$.
Let us now call $\chi_{m}$ a regular grid of $m$ nodes on $\Omega$, and call $\mathcal{F}_{m}^{c_{0},h}$ the family of FNOs trained on resolution $m$ with $c_{0}$ input channels and $h$ output channels of the dimension lifting operator. Thus, given a function $u\in\mathcal{H}_{1}$, we can express the restrictions $u|_{\chi_{m}}$ and $\mathcal{C}(u)|_{{\chi}_{m}}$ as vectors of size $md_{1}$ and $md_{2}$ respectively. This is more commonly arranged as a suitably shaped tensor, but we consider it a vector for convenience here. Let us also define $i_{m,m^{\prime}}:\mathcal{R}^{md_{1}}\to\mathcal{R}^{m^{\prime}d_{1}}$ and $\tilde{i}_{m,m^{\prime}}:\mathcal{R}^{md_{2}}\to\mathcal{R}^{m^{\prime}d_{2}}$ the up/downsampling operators for inputs and outputs, respectively. It is clear that, depending on the input $u\in\mathcal{H}_{1}$ one considers, the reconstruction gap
@@
can vary considerably, and the same applies to $\mathcal{C}(u)\in\mathcal{H}_{2}$. To simplify the notation, let us define $U_{m}=u|_{\chi_{m}}$ and $C_{m}(u)=\mathcal{C}(u)|_{{\chi}_{m}}$, where $C_{m}:\mathcal{H}_{1}\to\mathcal{R}^{d_{2}m}$. We also introduce the sets
$\mathcal{U}_{m}:=\left\{V\in\mathcal{R}^{md_{1}}:\,\,V=v|_{\chi_{m}}\text{ fora }v\in\mathcal{H}_{1}\right\},$
$\mathcal{V}_{m}:=\left\{V\in\mathcal{R}^{md_{2}}:V=\mathcal{C}(v)|_{\chi_{m}}\text{ for a }v\in\mathcal{H}_{1}\right\},$
@@
By the universal approximation property of FNOs, see [13], it is clear that, up to changing the considered FNO, one could make the associated error $\varepsilon_{m}(u)$ arbitrarily small. However, in this note, we keep it as a function and we do not focus on having necessarily small values. The FNO $\mathcal{N}$ can be found by training at the specific resolution determined by the grid $\chi_{m}$ defined over $\Omega\subset\mathcal{R}^{n}$. Suppose that when we test this FNO $\mathcal{N}\in\mathcal{F}_{m}^{c_{0},h}$ on inputs sampled at a different grid $\chi_{m^{\prime}}$ we have
@@
$\begin{split}\mathcal{R}_{1}:=\Big{\{}\tilde{i}_{m,m^{\prime}}\circ\mathcal{P}\circ\underline{\text{CAT}}\circ\text{CNN}_{\theta}\circ i_{m^{\prime},m}:\\ \mathcal{R}^{m^{\prime}d_{1}}\to\mathcal{R}^{m^{\prime}d_{2}}\,\big{|}\,\mathcal{P}\in\mathcal{F}_{m}^{c,h},\,\,\theta\in\mathcal{R}^{p}\Big{\}},\end{split}$ (6)
where $c=c_{0}+k$, $c_{0}$ is the number of input channels of our data points, and $k$ is the number of output channels of the CNN. We remark that the CNN depends on a set of trainable parameters collected in $\theta\in\mathcal{R}^{p}$, and $\underline{\text{CAT}}$ is the concatenation operation which concatenates the resized input $i_{m^{\prime},m}(u|_{\chi_{m^{\prime}}})\in\mathcal{U}_{m}$ with the CNN output. Furthermore, the notation CNN assumes the input to be reshaped into a suitable tensor, whereas here we work with vectors for convenience.
@@
$\begin{split}\mathcal{R}_{2}:=\Big{\{}\mathcal{P}\circ\text{CAT}\circ i_{m,m^{\prime}}\circ\text{CNN}_{\theta}\circ i_{m^{\prime},m}:\\ \mathcal{R}^{m^{\prime}d_{1}}\to\mathcal{R}^{m^{\prime}d_{2}}\,\big{|}\,\mathcal{P}\in\mathcal{F}_{m}^{c,h},\,\,\theta\in\mathcal{R}^{p}\Big{\}},\end{split}$ (7)
@@ 6 Conclusion
$\begin{split}\inf_{\text{Conv-FNO}\in\mathcal{R}_{2}}\left\|\text{Conv-FNO}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2}&=\inf_{\theta\in\mathcal{R}^{p}}\inf_{\mathcal{P}\in\mathcal{F}_{m}^{c,h}}\left\|\mathcal{P}\circ\text{CAT}\circ i_{m,m^{\prime}}\circ\text{CNN}_{\theta}\circ i_{m^{\prime},m}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2}\\ &\leq\inf_{\mathcal{P}\in\mathcal{F}_{m}^{c,h}}\left\|\mathcal{P}\circ\text{CAT}\circ i_{m,m^{\prime}}\circ\text{CNN}_{\hat{\theta}}\circ i_{m^{\prime},m}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2}\\ &=\inf_{\mathcal{N}\in\mathcal{F}_{m}^{c_{0},h}}\left\|\mathcal{N}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2},\end{split}$

### v2

@@ Abstract
Partial Differential Equation (PDE) problems often exhibit strong local spatial structures, and effectively capturing these structures is critical for approximating their solutions. Recently, the Fourier Neural Operator (FNO) has emerged as an efficient approach for solving these PDE problems. By using parametrization in the frequency domain, FNOs can efficiently capture global patterns. However, this approach inherently overlooks the critical role of local spatial features, as frequency-domain parameterized convolutions primarily emphasize global interactions without encoding comprehensive localized spatial dependencies. Although several studies have attempted to address this limitation, their extracted Local Spatial Features (LSFs) remain insufficient, and computational efficiency is often compromised. To address this limitation, we introduce a convolutional neural network (CNN)-based feature pre-extractor to capture LSFs directly from input data, resulting in a hybrid architecture termed Conv-FNO. Furthermore, we introduce two novel resizing schemes to make our Conv-FNO resolution invariant. In this work, we focus on demonstrating the effectiveness of incorporating LSFs into FNOs by conducting both a theoretical analysis and extensive numerical experiments. Our findings show that this simple yet impactful modification enhances the representational capacity of FNOs and significantly improves performance on challenging PDE benchmarks.
@@ 1 Introduction
Inspired by the success of CNO, we propose a simple yet powerful enhancement to the FNO architecture: the Convolutional Fourier Neural Operator (Conv-FNO). Specifically, our approach integrates a convolutional neural network (CNN) as a feature pre-extractor to extract LSFs directly from the input data before applying the Fourier transform. This preprocessing step effectively bridges the gap between local and global representations, allowing the model to retain fine-grained spatial details while leveraging the Fourier transform for efficient global interactions. Unlike previous methods focusing on frequency-domain modifications or downstream architectural enhancements, our approach explicitly targets input-stage feature extraction to ensure that critical local information is preserved from the outset.
Our proposed Conv-FNO preserves the resolution-invariance property of the original FNO through carefully designed resizing schemes, making it adaptable to grids of varying resolutions. Additionally, we introduce a specific variant, UNet-FNO, which incorporates a UNet-based encoder-decoder structure [26, 19] as the feature pre-extractor. This design leverages the hierarchical feature extraction capabilities of UNet to capture multi-scale spatial information, significantly enhancing the model’s ability to handle complex PDE solutions requiring both local and global representations.
@@ 4.2 Theoretical Analysis
The goal of an FNO is to approximate an operator between two Hilbert spaces $\mathcal{H}_{1},\mathcal{H}_{2}$, which we call $\mathcal{C}:\mathcal{H}_{1}\to\mathcal{H}_{2}$. We now suppose that the Hilbert space $\mathcal{H}_{1}$ is made of functions from $\Omega\subset\mathbb{R}^{n}$ and taking outputs in $\mathbb{R}^{d_{1}}$. The Hilbert space $\mathcal{H}_{2}$ takes instead inputs in ${\Omega}\subset\mathbb{R}^{n}$ and returns outputs in $\mathbb{R}^{d_{2}}$. This means that for an $x\in\Omega$, $u\in\mathcal{H}_{1}$, we have $u(x)\in\mathbb{R}^{d_{1}}$ and $\mathcal{C}(u)(x)\in\mathbb{R}^{d_{2}}$.
Let us now call $\chi_{m}$ a regular grid of $m$ nodes on $\Omega$, and call $\mathcal{F}_{m}^{c_{0},h}$ the family of FNOs trained on resolution $m$ with $c_{0}$ input channels and $h$ output channels of the dimension lifting operator. Thus, given a function $u\in\mathcal{H}_{1}$, we can express the restrictions $u|_{\chi_{m}}$ and $\mathcal{C}(u)|_{{\chi}_{m}}$ as vectors of size $md_{1}$ and $md_{2}$ respectively. This is more commonly arranged as a suitably shaped tensor, but we consider it a vector for convenience here. Let us also define $i_{m,m^{\prime}}:\mathbb{R}^{md_{1}}\to\mathbb{R}^{m^{\prime}d_{1}}$ and $\tilde{i}_{m,m^{\prime}}:\mathbb{R}^{md_{2}}\to\mathbb{R}^{m^{\prime}d_{2}}$ the up/downsampling operators for inputs and outputs, respectively. It is clear that, depending on the input $u\in\mathcal{H}_{1}$ one considers, the reconstruction gap
@@
can vary considerably, and the same applies to $\mathcal{C}(u)\in\mathcal{H}_{2}$. To simplify the notation, let us define $U_{m}=u|_{\chi_{m}}$ and $C_{m}(u)=\mathcal{C}(u)|_{{\chi}_{m}}$, where $C_{m}:\mathcal{H}_{1}\to\mathbb{R}^{d_{2}m}$. We also introduce the sets
$\mathcal{U}_{m}:=\left\{V\in\mathbb{R}^{md_{1}}:\,\,V=v|_{\chi_{m}}\text{ for a }v\in\mathcal{H}_{1}\right\},$
$\mathcal{V}_{m}:=\left\{V\in\mathbb{R}^{md_{2}}:V=\mathcal{C}(v)|_{\chi_{m}}\text{ for a }v\in\mathcal{H}_{1}\right\},$
@@
By the universal approximation property of FNOs, see [13], it is clear that, up to changing the considered FNO, one could make the associated error $\varepsilon_{m}(u)$ arbitrarily small. However, in this note, we keep it as a function and we do not focus on having necessarily small values. The FNO $\mathcal{N}$ can be found by training at the specific resolution determined by the grid $\chi_{m}$ defined over $\Omega\subset\mathbb{R}^{n}$. Suppose that when we test this FNO $\mathcal{N}\in\mathcal{F}_{m}^{c_{0},h}$ on inputs sampled at a different grid $\chi_{m^{\prime}}$ we have
@@
$\begin{split}\mathcal{R}_{1}:=\Big{\{}\tilde{i}_{m,m^{\prime}}\circ\mathcal{P}\circ\underline{\text{CAT}}\circ\text{CNN}_{\theta}\circ i_{m^{\prime},m}:\\ \mathbb{R}^{m^{\prime}d_{1}}\to\mathbb{R}^{m^{\prime}d_{2}}\,\big{|}\,\mathcal{P}\in\mathcal{F}_{m}^{c,h},\,\,\theta\in\mathbb{R}^{p}\Big{\}},\end{split}$ (6)
where $c=c_{0}+k$, $c_{0}$ is the number of input channels of our data points, and $k$ is the number of output channels of the CNN. We remark that the CNN depends on a set of trainable parameters collected in $\theta\in\mathbb{R}^{p}$, and $\underline{\text{CAT}}$ is the concatenation operation which concatenates the resized input $i_{m^{\prime},m}(u|_{\chi_{m^{\prime}}})\in\mathcal{U}_{m}$ with the CNN output. Furthermore, the notation CNN assumes the input to be reshaped into a suitable tensor, whereas here we work with vectors for convenience.
@@
$\begin{split}\mathcal{R}_{2}:=\Big{\{}\mathcal{P}\circ\text{CAT}\circ i_{m,m^{\prime}}\circ\text{CNN}_{\theta}\circ i_{m^{\prime},m}:\\ \mathbb{R}^{m^{\prime}d_{1}}\to\mathbb{R}^{m^{\prime}d_{2}}\,\big{|}\,\mathcal{P}\in\mathcal{F}_{m}^{c,h},\,\,\theta\in\mathbb{R}^{p}\Big{\}},\end{split}$ (7)
@@ 6 Conclusion
$\begin{split}\inf_{\text{Conv-FNO}\in\mathcal{R}_{2}}\left\|\text{Conv-FNO}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2}&=\inf_{\theta\in\mathbb{R}^{p}}\inf_{\mathcal{P}\in\mathcal{F}_{m}^{c,h}}\left\|\mathcal{P}\circ\text{CAT}\circ i_{m,m^{\prime}}\circ\text{CNN}_{\theta}\circ i_{m^{\prime},m}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2}\\ &\leq\inf_{\mathcal{P}\in\mathcal{F}_{m}^{c,h}}\left\|\mathcal{P}\circ\text{CAT}\circ i_{m,m^{\prime}}\circ\text{CNN}_{\hat{\theta}}\circ i_{m^{\prime},m}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2}\\ &=\inf_{\mathcal{N}\in\mathcal{F}_{m}^{c_{0},h}}\left\|\mathcal{N}(u|_{\chi_{m^{\prime}}})-C_{m^{\prime}}(u)\right\|_{2},\end{split}$